A Population Has A Mean Of 73.5 And A Standard Deviation Of 2.5


1
Expert's answer
2022-02-24T05:43:20-0500

Let "X_1, X_2, ..., X_n" be a random sample from a distribution with mean "\mu" and variance "\sigma^2". Then by the Central Limit Theorem if "n" is sufficiently large, "\bar{X}" has approximately a normal distribution with "\mu_{\bar{X}}=\mu" and "\sigma^2_{\bar{X}}=\sigma^2/n."

The Central Limit Theorem can generally be used if "n>30."

A.

Given "\mu=73.5, \sigma=2.5, n=30."

Assume "\bar{X}\sim N(\mu, \sigma^2/n)"

"\mu_{\bar{X}}=\mu=73.5"

"\sigma_{\bar{X}}=\sigma/\sqrt{n}=2.5/\sqrt{30}\approx0.4564"

B.


"P(\bar{X}<72)=P(Z<\dfrac{72-73.5}{2.5/\sqrt{30}})"

"\approx P(Z<-3.2863)\approx0.0005"